# Accelerated Ostwald ripening by chemical activity

### v1

@@ Acknowledgements
In the main text, where we have taken for simplicity $\ell_{\mathrm{l}}=\ell_{\mathrm{g}}\equiv\ell$, Eq. (21) contains $g(x)=G(x,1)$.
SII.5 Droplet growth law
Equipped with the gas phase flux, Eq. (S49) (where the linearization is adequate since the mixture is only slightly supersaturated, $0<\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}\ll 1$), and the liquid-phase flux, Eq. (S72) (which we have obtained without linearization, though we needed to assume that the reactions do not destablize the dense phase), we proceed to compute the radius growth rate, expressed from either the sticky and nonsticky fluxes. From these equations, we will extract $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ and the sought-after $\mathrm{d}R/\mathrm{d}t$.
SII.5.1 Droplet growth in terms of particle flux
First, we derive Eqs. (5) and (12) of the main text. We integrate Eq. (S37) (assuming spherical symmetry) over the entire container,
where in the last equality we recalled the fast-surface-thermodynamics boundary condition, $\bar{\phi}_{\alpha}(R,t)=\bar{\phi}_{\alpha}^{\mathrm{L}}(R)$ ($\alpha=\mathrm{l},\mathrm{g}$), and we brought back the radius dependencies of $\bar{\phi}_{\alpha}^{\mathrm{L}}(R)$ for clarity. Upon comparing Eqs. (S74) and (S75), we cancel out the reaction term $\bar{\upsilon}\int_{0}^{\infty}\mathrm{d}rr^{2}s(\bar{\phi}(r,t))$ and isolate $\mathrm{d}R(t)/\mathrm{d}t$
$[\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}(R(t))-\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}(R(t))]\frac{\mathrm{d}R(t)}{\mathrm{d}t}=\bar{J}_{\mathrm{l}}(R(t))-\bar{J}_{\mathrm{g}}(R(t)),$ (S76)
which yields Eqs. (5) and (12) of the main text.
@@
$[\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}(R(t))-\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}(R(t))]\frac{\mathrm{d}R(t)}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R(t)}[\bar{\varphi}^{\infty}-\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}(R(t))]+\bar{\upsilon}\ell_{\mathrm{g}}\left[s_{\mathrm{g}}+s_{\mathrm{l}}G\left(\frac{R(t)}{\ell_{\mathrm{g}}},\frac{\ell_{\mathrm{l}}}{\ell_{\mathrm{g}}}\right)\right].$ (S77)
@@
which is the generalization of Eq. (21) of the main text to non-first-order reaction mechanisms in the gas phase, non-one-to-one stoichiometric ratios, and $\ell_{\mathrm{l}}\neq\ell_{\mathrm{g}}$. To obtain Eq. (21), based on Sec. SII.3.2, we define $\bar{k}_{\mathrm{g}}=(k_{\mathrm{s}\to\mathrm{ns}},-k_{\mathrm{ns}\to\mathrm{s}})^{\mathrm{T}}$ (which are constant rates in the linearization of Sec. SII.3.2), from which $\ell_{\mathrm{g}}=[D_{\mathrm{g}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})]^{1/2}$, $s_{\mathrm{g}}=\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, and assume that $\ell_{\mathrm{l}}=\ell_{\mathrm{g}}\equiv\ell$.
More generally, since $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}\to 0$ is subdominant and $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}\to 1$ is determined independently of $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, neither $s_{\mathrm{l}}$ nor $\ell_{\mathrm{l}}$ depend on $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$. Therefore, only the total zeroth-order reaction rate $s_{\mathrm{g}}=s(\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})$ and inter-reaction distance $\ell_{\mathrm{g}}=\{D_{\mathrm{g}}/[-\bar{k}_{\mathrm{g}}^{\mathrm{T}}(\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})\bar{\upsilon}]\}^{1/2}$ might depend on $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$. (In the simpler case of the main text and Sec. SII.3.2, $\ell_{\mathrm{g}}$ does not after all.) Upon determining $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ from Eq. (S79), the radius growth law can be found explicitly from Eq. (S78). We proceed to do so in the different limits of $R$ versus $\ell_{\mathrm{g}}$ and $\ell_{\mathrm{l}}$.
SII.5.2 Small droplet
@@
Indeed, $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}\simeq\varphi_{\mathrm{ns}}^{\infty}$, thus justifying the linearization ansatz (Eq. (S38)) for the nonsticky species as well. (To recover Eq. (23) of the main text from this general expression, insert $s_{\mathrm{g}}=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, $\ell_{\mathrm{g}}=[D_{\mathrm{g}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})]^{1/2}$, and $\varphi_{\mathrm{ns}}^{\infty}=k_{\mathrm{s}\to\mathrm{ns}}\varphi^{\infty}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})$.)
@@
(To obtain Eq. (24) of the main text, insert the above conditions along with $|\upsilon_{\mathrm{s}}|=1$.) The first term corresponds to passive Ostwald ripening, the second term is a subdominant acceleration due to reactions in the dilute phase, and the third term stems from the growth-opposing reactions in the dense droplet.
SII.5.3 Large droplet
@@
Here, without further assumptions (as in Sec. SII.3.2 and in the main text) $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}=\phi_{\mathrm{ns}\mathrm{g}}^{\circ}+\lambda_{\mathrm{ns}\mathrm{g}}/R$ can only be found implicitly. To leading-order in $\ell_{\mathrm{g}}/R,\lambda_{\mathrm{ns}\mathrm{g}}/R\ll 1$, $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}$ must satisfy implicitly
$-s_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})\ell_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})=s_{\mathrm{l}}\ell_{\mathrm{l}}.$ (S84)
That is, since the reactions occur with rates proportional to the droplet surface, while the droplet volume can only grow at a rate proportional to the radius due to the mass conservation constraint (Eq. (S78)), the reactions must balance each other to leading order for a large droplet, $(4\pi R^{2}\ell_{\mathrm{g}})s_{\mathrm{g}}+(4\pi R^{2}\ell_{\mathrm{l}})s_{\mathrm{l}}=0$, where the corresponding reactive shell widths around the interface are $\ell_{\alpha}$ ($\alpha=\mathrm{l},\mathrm{g}$). (In the limit considered in Sec. SII.3.2 and the main text, $s_{\mathrm{g}}=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ and $\ell_{\mathrm{g}}=\ell_{\mathrm{l}}$, Eq. (S84) rearranges as $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}=(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}+s_{\mathrm{l}}/k_{\mathrm{ns}\to\mathrm{s}}$, which is the first line of Eq. (25) of the main text.) Since, to avoid size control, $s_{\mathrm{l}}<(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})[\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{L}]$, both $\mathrm{s}_{\mathrm{l}}$ and $s_{\mathrm{g}}$ are of order $\varphi^{\infty}$, and hence the expression for $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}$ will be of order $\varphi^{\infty}$; thus, the linearization (Eq. (S38)) is justified in this limit as well.
@@
(In the case considered in the main text, where $D_{\mathrm{g}}/\ell_{\mathrm{g}}^{2}=k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}$, $k_{\mathrm{ns}}=-k_{\mathrm{ns}\to\mathrm{s}}$, $\ell_{l}=\ell_{\mathrm{g}}=\ell$, and the above expression for $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}=(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})\phi_{\mathrm{s}\mathrm{g}}^{\circ}+s_{\mathrm{l}}/k_{\mathrm{ns}\to\mathrm{s}}$, Eq. (S85) coincides with the second row of Eq. (25) of the main text. Inserting these expressions for $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}=\phi_{\mathrm{ns}\mathrm{g}}^{\circ}+\lambda_{\mathrm{ns}\mathrm{g}}/R$ in Eq. (S78) yields Eq. (26) of the main text; without these additional assumptions allowing us to express $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ explicitly, we cannot immediately conclude that acceleration occurs.) Should $\ell_{\mathrm{g}}$ also depend on $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, one would need to expand it in small $\lambda_{\mathrm{ns}\mathrm{g}}/R$ in Eq. (S83), too.
SII.5.4 Intermediate droplet
What we mean to explore in this section, for completeness, is a situation where reactions occur everywhere throughout the droplet, even for larger droplets, but in the regime where $R\ll\ell_{\mathrm{l}}$, so that the volume fraction of the particles remains constant throughout the droplet, $\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}$. (For instance, this might be the case in biocondensates, where the enzymes might be attached to the sticky species and thus always present and active inside the dense phase.) The limit where $R\ll\ell_{\mathrm{g}}$ was explored in Sec. SII.5.2. Here, we aim to explore the intermediate regime, $\ell_{\mathrm{g}}\ll R\ll\ell_{\mathrm{l}}$. Therefore, Eq. (S80) is still valid. However to avoid size control (to be able to cross from $R\ll\ell_{\mathrm{g}}$ to $R\gg\ell_{\mathrm{g}}$), we must assume that the total reaction rate inside the droplet is much smaller than in the surrounding gas phase, $s_{\mathrm{l}}R^{3}\ll|s_{\mathrm{g}}|R^{2}\ell_{\mathrm{g}}$; otherwise the last term in Eq. (S80) will strongly drive the shrinkage of the droplet.
As in Sec. SII.5.3, $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ is only found implicitly. With the assumption $s_{\mathrm{l}}\ll|s_{\mathrm{g}}|\ell_{\mathrm{g}}/R$, the criterion reads
$s_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})=0;$ (S86)
cf. Eq. (S84) with $s_{\mathrm{l}}\to 0$ (and $\ell_{\mathrm{g}}\neq 0$). Since in this combination of limits, reactions only create material from the gas-phase side of the interface, this equation states that the gas-phase reactions must balance out on their own. Next, we again find $\lambda_{\mathrm{ns}\mathrm{g}}$ in the case where $\ell_{\mathrm{g}}=\mathrm{const}$,
@@
$\frac{\lambda_{\mathrm{ns}\mathrm{g}}}{R}=\frac{D_{\mathrm{g}}}{-k_{\mathrm{ns}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})\ell_{\mathrm{g}}^{2}}\left(\varphi^{\infty}_{\mathrm{ns}}-\phi_{\mathrm{ns}\mathrm{g}}^{\circ}\right)\frac{\ell_{\mathrm{g}}}{R}+\frac{s_{\mathrm{l}}}{3[-k_{\mathrm{ns}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})]}\frac{R}{\ell_{\mathrm{g}}}.$ (S87)
Thus, it is as if there are no reactions inside the droplet (cf. Eq. (S85) with $s_{\mathrm{l}}\to 0$ as the first term in Eq. (S87)), to which we add retroactively the small correction due to the subdominant dense-phase reactions.
Explicitly, in the case considered in the main text and Sec. SII.3.2, we find
@@
$\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}=\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\left(1+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}}\right)(\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ})\frac{\ell_{\mathrm{g}}}{R}+\frac{s_{\mathrm{l}}}{3k_{\mathrm{ns}\to\mathrm{s}}}\frac{R}{\ell_{\mathrm{g}}}.$ (S88)
Inserting this in Eq. (S78) and taking the limit of late ripening (so $(\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ})(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})\ell_{\mathrm{g}}\ll\lambda$), we find

### v2

@@ Acknowledgements
This is Eq. (7) of the main text, which therefore avoids the unjustified dense-phase linearization. Instead, we have compromized for “phenomenological” dense-phase rate constant $s_{\mathrm{l}}$ and reactive equilibration distance $\ell_{\mathrm{l}}$.
SII.6 Droplet growth law
Equipped with the gas phase flux, Eq. (S53) (where the linearization is adequate since the mixture is only slightly supersaturated, $0<\phi_{\mathrm{s}\mathrm{g}}^{\circ}\lesssim\varphi_{\mathrm{s}}^{\infty}\ll 1$), and the liquid-phase flux, Eq. (S76) (which we have obtained without linearization, though we needed to assume that the reactions do not destablize the dense phase), we proceed to compute the radius growth rate, expressed from either the sticky and nonsticky fluxes. From these equations, we will extract $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ and the sought-after $\mathrm{d}R/\mathrm{d}t$.
@@
$\bar{\phi}_{\mathrm{l}}^{\mathrm{L}}(R(t))-\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}(R(t))]\frac{\mathrm{d}R(t)}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R(t)}[\bar{\varphi}^{\infty}-\bar{\phi}_{\mathrm{g}}^{\mathrm{L}}(R(t))]+\bar{\upsilon}\ell_{\mathrm{g}}\left[s_{\mathrm{g}}+s_{\mathrm{l}}G\left(\frac{R(t)}{\ell_{\mathrm{g}}},\frac{\ell_{\mathrm{l}}}{\ell_{\mathrm{g}}}\right)\right].$ (S78)
@@
from which $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}(R)$ will be extracted. Since $\phi_{\mathrm{ns}\mathrm{l}}^{\mathrm{L}}\to 0$ is subdominant and $\phi_{\mathrm{s}\mathrm{l}}^{\mathrm{L}}\to 1$ is determined independently of $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$, neither $s_{\mathrm{l}}$ nor $\ell_{\mathrm{l}}$ depend on $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$. Therefore, only the total zeroth-order reaction rate $s_{\mathrm{g}}=s(\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})$ and inter-reaction distance $\ell_{\mathrm{g}}=\{D_{\mathrm{g}}/[-\bar{k}_{\mathrm{g}}^{\mathrm{T}}(\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}},\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}})\bar{\upsilon}]\}^{1/2}$ might depend on $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$. (In the simpler case of the main text and Sec. SII.4.2, $\ell_{\mathrm{g}}$ does not after all.) Upon determining $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}$ from Eq. (S80), the radius growth law can be found explicitly from Eq. (S79). We proceed to do so in the different limits of $R$ versus $\ell_{\mathrm{g}}$ and $\ell_{\mathrm{l}}$. The physical outcomes obtained for various values of $s_{\mathrm{l}}$ and droplet size $R$ are sketched in Fig. S1.
Figure S1: “Phase diagram” of droplet dynamics as a function of total dense phase reaction rate $s_{\mathrm{l}}$ and droplet size $R(t)$. Reactions in the dilute phase may accelerate Ostwald ripening (cf. Fig. 2 of the main text), whereas reactions in the dense phase hinder droplet growth and ripening by converting particles from sticky to nonsticky. Small droplets with $R\ll\lambda_{\mathrm{s}\mathrm{g}}/(\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ})$ always tend to shrink due to large Laplace pressure. Other regions are approximately bounded by the inter-reaction diffusion length $\ell_{\mathrm{g}}$. For very fast dense phase reactions, $s_{\mathrm{l}}\ell_{\mathrm{l}}^{2}/D_{\mathrm{g}}\gg\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}$, size control is inevitable with stable radius $R_{\mathrm{sc}}^{2}=3D_{\mathrm{g}}(\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ})/s_{\mathrm{l}}$. Otherwise, droplets can reach the regime $R\gg\ell_{\mathrm{l}},\ell_{\mathrm{g}}$, where: For intermediate dense phase reaction rate, $(1+\ell_{\mathrm{g}}/\ell_{\mathrm{l}})\lambda_{\mathrm{s}\mathrm{g}}/\ell_{\mathrm{l}}\ll s_{\mathrm{l}}\ell_{\mathrm{l}}^{2}/D_{\mathrm{g}}\ll\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}$, a negative effective surface tension prevents small droplets from shrinking. For slow dense phase reaction rate, $s_{\mathrm{l}}\ell_{\mathrm{l}}^{2}/D_{\mathrm{g}}\ll(1+\ell_{\mathrm{g}}/\ell_{\mathrm{l}})\lambda_{\mathrm{s}\mathrm{g}}/\ell_{\mathrm{l}}$, Ostwald ripening occurs with an accelerated rate. Parameters: $R$, droplet radius; $s_{\mathrm{l}}$, total dense phase reaction rate; $\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}$, dilute phase volume fraction relative to the coexistence value; $\lambda_{\mathrm{s}\mathrm{g}}$, Laplace-pressure correction to the volume fraction of the sticky species in the gas phase; $\ell_{\mathrm{l}},\ell_{\mathrm{g}}$, characteristic distance between reactions in each phase; $D_{\mathrm{g}}$, dilute-phase diffusion constant.
SII.6.1 Small droplet
@@
Indeed, $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}\simeq\varphi_{\mathrm{ns}}^{\infty}$, thus justifying the linearization ansatz (Eq. (S42)) for the nonsticky species as well.
@@
The first term corresponds to passive Ostwald ripening, the second term is the acceleration due to reactions in the dilute phase, and the third term stems from the growth-opposing reactions in the dense droplet. We analyze size control by rewriting Eq. (S83) as
$\frac{\mathrm{d}R(t)}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R(t)}\left\{[\varphi^{\infty}_{\mathrm{s}}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R(t))]+|\upsilon_{\mathrm{s}}|\frac{-s_{\mathrm{g}}(\varphi^{\infty}_{\mathrm{ns}})}{-\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\upsilon}}\frac{R(t)}{\ell_{\mathrm{g}}(\varphi^{\infty}_{\mathrm{ns}})}-\frac{|\upsilon_{\mathrm{s}}|s_{\mathrm{l}}R^{2}(t)}{3D_{\mathrm{g}}}\right\},$ (S84)
Since no reactions occur in the absence of material, $s(\bar{0})=0$, we can expect at least $s_{\mathrm{g}}]/[-\bar{k}_{\mathrm{g}}^{\mathrm{T}}\bar{\upsilon}]=\mathcal{O}(\varphi)$, meaning that the acceleration due to gas-phase reactions is subleading — the second term is smaller than the first one by $R/\ell_{\mathrm{g}}\ll 1$. The last term, originating from dense-phase reactions, competes with the passive diffusive growth, and it can prevent it entirely once all droplets reach the stable size-control radius
$R_{\mathrm{sc}}=\sqrt{\frac{3D_{\mathrm{g}}}{|\upsilon_{\mathrm{s}}|s_{\mathrm{l}}}[\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(R_{\mathrm{sc}})]}.$ (S85)
Note that as droplets grow, the supersaturated dilute phase loses material, meaning that $\varphi_{\mathrm{s}}^{\infty}$ is indirectly also affected by $R_{\mathrm{sc}}$. Thus, to avoid size control, we need to transition to the next regime where the dense-phase flux reduces to $R^{2}J_{\mathrm{l}}\sim s_{\mathrm{l}}R^{2}$ (for $R\gg\ell_{\mathrm{l}}$) prior to reaching $R_{\mathrm{sc}}$. This translates to the requirement $R_{\mathrm{sc}}\gg\ell_{\mathrm{l}}$, or
$\frac{\ell_{\mathrm{l}}^{2}s_{\mathrm{l}}}{D_{\mathrm{g}}}\ll\varphi_{\mathrm{s}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\mathrm{L}}(\ell_{\mathrm{l}}),$ (S86)
shown in the main text.
SII.6.2 Large droplet
@@
Here, without further assumptions (as in Sec. SII.4.2 and in the main text) $\phi_{\mathrm{ns}\mathrm{g}}^{\mathrm{L}}=\phi_{\mathrm{ns}\mathrm{g}}^{\circ}+\lambda_{\mathrm{ns}\mathrm{g}}/R$ can only be found implicitly. To leading-order in $\ell_{\mathrm{g}}/R,\lambda_{\mathrm{ns}\mathrm{g}}/R\ll 1$, $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}$ must satisfy implicitly
$s_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})\ell_{\mathrm{g}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})=s_{\mathrm{l}}\ell_{\mathrm{l}}.$ (S88)
That is, since the reactions occur with rates proportional to the droplet surface, while the droplet volume can only grow at a rate proportional to the radius due to the mass conservation constraint (Eq. (S79)), the reactions must balance each other to leading order for a large droplet, $(4\pi R^{2}\ell_{\mathrm{g}})s_{\mathrm{g}}+(4\pi R^{2}\ell_{\mathrm{l}})s_{\mathrm{l}}=0$, where the corresponding reactive shell widths around the interface are $\ell_{\alpha}$ ($\alpha=\mathrm{l},\mathrm{g}$).
@@
Combining, we find
@@
$\frac{\mathrm{d}R(t)}{\mathrm{d}t}=\frac{D_{\mathrm{g}}}{R(t)}\left[(\varphi^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ}-|\upsilon_{\mathrm{s}}|\phi_{\mathrm{ns}\mathrm{g}}^{\circ})-\frac{\lambda_{\mathrm{s}\mathrm{g}}}{R(t)}\left(1-\frac{s_{\mathrm{l}}\ell_{\mathrm{l}}^{2}}{[-k_{\mathrm{ns}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})]\ell_{\mathrm{g}}^{2}}\frac{\ell_{\mathrm{g}}}{\lambda_{\mathrm{s}\mathrm{g}}}+\frac{D_{\mathrm{g}}(\varphi_{\mathrm{ns}}^{\infty}-\phi_{\mathrm{ns}\mathrm{g}}^{\circ})}{[-k_{\mathrm{ns}}(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})]\ell_{\mathrm{g}}\lambda_{\mathrm{s}\mathrm{g}}}\right)\right].$ (S90)
To make progress, we consider the gas-phase reaction mechanism considered in the text, $s(\bar{\phi}_{\mathrm{g}})=k_{\mathrm{s}\to\mathrm{ns}}\phi_{\mathrm{s}\mathrm{g}}-k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}$ with $\upsilon_{\mathrm{s}}=-1$. In this case, $s(\phi_{\mathrm{ns}\mathrm{g}}^{\circ})=k_{\mathrm{ns}\to\mathrm{s}}\phi_{\mathrm{ns}\mathrm{g}}^{\circ}-k_{\mathrm{s}\to\mathrm{ns}}(\phi_{\mathrm{s}\mathrm{g}}^{\circ}+\lambda_{\mathrm{s}\mathrm{g}}/R)$, $\ell_{\mathrm{g}}^{2}=D_{\mathrm{g}}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})$, and $k_{\mathrm{ns}}=k_{\mathrm{ns}\to\mathrm{s}}$. Consequently, we have $\phi_{\mathrm{ns}\mathrm{g}}^{\circ}=(k_{\mathrm{s}\to\mathrm{ns}}/k_{\mathrm{ns}\to\mathrm{s}})(\phi_{\mathrm{s}\mathrm{g}}^{\circ}+\lambda_{\mathrm{s}\mathrm{g}}/R)+(s_{\mathrm{l}}/k_{\mathrm{ns}\to\mathrm{s}})(\ell_{\mathrm{l}}/\ell_{\mathrm{g}})$ and $\varphi_{\mathrm{ns}}^{\infty}=k_{\mathrm{s}\to\mathrm{ns}}\varphi^{\infty}/(k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}})$, so
@@
$\frac{\mathrm{d}R(t)}{\mathrm{d}t}=\frac{D_{\mathrm{eff}}}{R(t)}\left[(\varphi_{\mathrm{eff}}^{\infty}-\phi_{\mathrm{s}\mathrm{g}}^{\circ})-\frac{\lambda_{\mathrm{eff}}}{R(t)}\right],$ (S91)
where
$\frac{D_{\mathrm{eff}}}{D_{\mathrm{g}}}$ $=$ $1+\frac{k_{\mathrm{s}\to\mathrm{ns}}}{k_{\mathrm{ns}\to\mathrm{s}}},$ (S92)
$\varphi_{\mathrm{eff}}^{\infty}$ $=$ $\frac{k_{\mathrm{ns}\to\mathrm{s}}\varphi^{\infty}-s_{\mathrm{l}}(\ell_{\mathrm{l}}/\ell_{\mathrm{g}})}{k_{\mathrm{s}\to\mathrm{ns}}+k_{\mathrm{ns}\to\mathrm{s}}},$ (S93)